# Regularized Langevin Dynamics for Combinatorial Optimization

### v3

@@ 5.3 Ablation Study
To verify the effectiveness of the proposed regularization in RLD, we conduct the ablation study on RLSA and RLNN, respectively. We first compare RLD with the standard discrete LD (Zhang et al., 2022) for SA, searching the step size $\alpha$ over the set $\{0.1,0.01,0.001\}$. All other hyperparameters are kept the same as in RLSA. Figure 1 compares the dynamics of the primal gap (Berthold, 2014) across the sampling process. Here, the primal gap on each instance is defined as
@@
In fact, the idea of RLNN also parallels well‐known reinforcement learning techniques to encourage exploration, such as curiosity-driven exploration (Pathak et al., 2017) and soft policies (Haarnoja et al., 2017, 2018).
@@ A.2 Implementation of RLNN
RLNN is parameterized by a five-layer GCN (Kipf & Welling, 2017) with 128 hidden dimensions. A linear layer is first used to project the input $\mathbf{x}$ to a 128-dim embedding $\mathbf{H}^{0}$. Each layer of GCN performs the following update:
@@
In terms of training time, our model completes training on small-scale graphs (except ER-[700700800800]) in under half an hour using eight RTX A6000 GPUs, and the larger instances finish within one hour; on a server with ten RTX 2080 Ti GPUs, runtimes are slightly longer, but the longest experiment still finishes within three hours. By comparison, baseline methods such as DiffUCO (Sanokowski et al., 2024) require two days to train a single model on RB–[200200300300]. We attribute this efficiency advantage to the local training objective of RLNN, which avoids estimating long-term high-variance reward signals with standard reinforcement learning methods, as employed by DiffUCO.
@@ Appendix B Comparison under Longer Runtime
Appendix C Example Code for RLSA
@@
1def energy_func(A, b, x, penalty_coeff=1.02):
2 """
3␣␣␣␣Theenergyfunctionis:b^Tx+penalty_coeff*x^TAx
4␣␣␣␣Returntheenergyandthegradient
5␣␣␣␣"""
6
7 L = A@x
8 energy = torch.sum(x*(penalty_coeff*L+b), dim=0)
9 grad = 2*penalty_coeff*L+b
10
11 return energy, grad
12
13def RLSA(graph, tau0, step_size, num_runs, num_steps, penalty_coeff):
14 """
15␣␣␣␣graph:thegraphobjectintorch_geometric
16␣␣␣␣num_runs:thenumberofparallelSAprocesses
17␣␣␣␣num_steps:thenumberofSAsteps
18␣␣␣␣tau0:theinitialtemperature
19␣␣␣␣"""
20
21 # initialization
22 num_nodes = graph.num_nodes
23
24 A = torch.sparse_coo_tensor(
25 graph.edge_index,
26 graph.edge_weight,
27 torch.Size((num_nodes, num_nodes))
28 )
29 x = torch.randint(0,2, (num_nodes, num_runs))
30
31 energy, grad = energy_func(A, graph.b, x, penalty_coeff)
32 best_energy = energy
33 best_sol = x.clone()
34
35 # SA
36 for epoch in range(num_steps):
37 # annealing
38 tau = tau0*(1-epoch/num_steps)
39
40 # sampling
41 delta = grad*(2*x-1)/2
42 term2 = -torch.kthvalue(
43 -delta,
44 step_size,
45 dim=0,
46 keepdim=True
47 ).values
48
49 flip_prob = torch.sigmoid((delta-term2)/tau)
50 $r$ = torch.rand_like(x.data)
51 x = torch.where(rr<flip_prob, 1-x, x)
52
53 # update loss
54 energy, grad = energy_func(A, graph.b, x, penalty_coeff)
55 to_update = energy<best_energy
56 best_sol[:,to_update] = x[:,to_update]
57 best_energy[to_update] = energy[to_update]
58
59 return best_energy, best_sol

### v4

@@ 5.3 Ablation Study
To verify the effectiveness of the proposed regularization in RLD, we conduct the ablation study on RLSA and RLNN, respectively. We first compare RLD with the standard discrete LD (zhang2022langevinlike) for SA, searching the step size $\alpha$ over the set $\{0.1,0.01,0.001\}$. All other hyperparameters are kept the same as in RLSA. Figure 1 compares the dynamics of the primal gap (Berthold2014HeuristicAI) across the sampling process. Here, the primal gap on each instance is defined as
@@
In fact, the idea of RLNN also parallels well‐known reinforcement learning techniques to encourage exploration, such as curiosity-driven exploration (pathakICMl17curiosity) and soft policies (haarnoja2017reinforcement; pmlr-v80-haarnoja18b).
@@ A.2 Implementation of RLNN
RLNN is parameterized by a five-layer GCN (kipf2017semisupervised) with 128 hidden dimensions. A linear layer is first used to project the input $\mathbf{x}$ to a 128-dim embedding $\mathbf{H}^{0}$. Each layer of GCN performs the following update:
@@
In terms of training time, our model completes training on small-scale graphs (except ER-[700700800800]) in under half an hour using eight RTX A6000 GPUs, and the larger instances finish within one hour; on a server with ten RTX 2080 Ti GPUs, runtimes are slightly longer, but the longest experiment still finishes within three hours. By comparison, baseline methods such as DiffUCO (SanokowskiHL24) require two days to train a single model on RB–[200200300300]. We attribute this efficiency advantage to the local training objective of RLNN, which avoids estimating long-term high-variance reward signals with standard reinforcement learning methods, as employed by DiffUCO.
@@ Appendix B Comparison under Longer Runtime
Appendix C RLSA Pseudocode
@@
RLSA Pseudocode def energy_func(A, b, x, penalty_coeff=1.02): """ The energy function is: b^Tx+penalty_coeff*x^TAx Return the energy and the gradient """ L = A@x energy = torch.sum(x*(penalty_coeff*L+b), dim=0) grad = 2*penalty_coeff*L+b return energy, grad def RLSA(graph, tau0, step_size, num_runs, num_steps, penalty_coeff): """ graph: the graph object in torch_geometric num_runs: the number of parallel SA processes num_steps: the number of SA steps tau0: the initial temperature """ # initialization num_nodes = graph.num_nodes A = torch.sparse_coo_tensor( graph.edge_index, graph.edge_weight, torch.Size((num_nodes, num_nodes)) ) x = torch.randint(0,2, (num_nodes, num_runs)) energy, grad = energy_func(A, graph.b, x, penalty_coeff) best_energy = energy best_sol = x.clone() # SA for epoch in range(num_steps): # annealing tau = tau0*(1-epoch/num_steps) # sampling delta = grad*(2*x-1)/2 term2 = -torch.kthvalue( -delta, step_size, dim=0, keepdim=True ).values flip_prob = torch.sigmoid((delta-term2)/tau) $r$ = torch.rand_like(x.data) x = torch.where(rr<flip_prob, 1-x, x) # update loss energy, grad = energy_func(A, graph.b, x, penalty_coeff) to_update = energy<best_energy best_sol[:,to_update] = x[:,to_update] best_energy[to_update] = energy[to_update] return best_energy, best_sol